OpenStudy (anonymous):

3. What is the value of x? 5 2.5 7.5 10

OpenStudy (anonymous):

I tried solving for x but I got 3.25

rishavraj (rishavraj):

IT WONT LET ME ACCESS THE PAGE -_-

OpenStudy (anonymous):



OpenStudy (anonymous):

@rishavraj

rishavraj (rishavraj):

why 3.25 ?? apply the concept of similar triangles :))

OpenStudy (anonymous):

what is the concept of similar triangles, my teacher isnt the best for my learning style.

rishavraj (rishavraj):

which means \[\frac{ x }{ x+5 }~~=~~\frac{ x - 10 }{ x + 1 }\]

OpenStudy (anonymous):

so does that mean I would write it as

OpenStudy (anonymous):

I still dont understand do you think you can walk me through the steps?

rishavraj (rishavraj):

sorry for late reply....internet lagging :?

OpenStudy (anonymous):

nope your good

OpenStudy (anonymous):

however I just dont understand what we are multiplying because there are more than two different x equations

rishavraj (rishavraj):

see the equation I mentiond we need to solve it to get x

rishavraj (rishavraj):

lemme know if u hav been taught about similar triangles??

OpenStudy (anonymous):

yes i see we have to solve for x